### v1

@@ Abstract
Context. The Gaia mission discovered many new candidate $\beta$ Cephei ($\beta$ Cep) pulsators, whose variability classification is meanwhile confirmed from Transiting Exoplanet Survey Satellite (TESS) space photometry of the nominal mission.
Aims. We aim to analyse all currently available TESS data for these $\beta$ Cep pulsators, of which 145 were new discoveries, in order to exploit their asteroseismic potential. $\beta$ Cep stars belong to an under-represented class of pulsators in the current space photometry revolution while being of critical importance to improve evolution models of massive stars.
Methods. We extracted light curves for 216 star from the TESS full-frame images and performed pre-whitening to derive all significant frequencies. Based on Gaia Data Release 3, we deduced stellar properties and compared them to those of known $\beta$ Cep stars from the literature. We developed a methodology to identify the dominant pulsation modes of the $\beta$ Cep stars from Gaia and TESS amplitude ratios and from the detection of rotationally-split multiplets. We used grid modelling to gain insights into the population of $\beta$ Cep stars.
Results. With the combination of TESS and Gaia, we successfully identified the mode degrees for 176 stars in our sample. We find the majority to have a dominant dipole non-radial mode ($l=1$). Many non-radial modes show splittings in their TESS frequency spectra. This allowed us to assemble a large set of split multiplets in $\beta$ Cep stars and to calculate their envelope rotation, spin parameter, and the level of differential envelope-to-surface rotation. For the latter, we find an upper limit of 4, with most stars rotating almost rigidly. We also provide the asymmetries of the multiplets, which range from -0.15 to 0.15 with the majority being positive. Based on grid modelling, we provide mass, convective core mass, and age distributions for 164 stars.
Conclusions. By combining Gaia DR3 and TESS, we enable asteroseismology of $\beta$ Cep pulsators as a population. Our study prepares for future detailed modelling based on individual frequencies of identified modes leading towards a better understanding of these massive pulsators, as crucial probes of stellar evolution theory.
@@ 1 Introduction
$\beta$ Cephei ($\beta$ Cep) stars are core-hydrogen burning massive stars ($8\lesssim M/M_{\sun}\lesssim 25$) pulsating in lower-order pressure (p-) or gravity (g-) modes. Despite being known for more than a century (Frost, 1906; Guthnick, 1913) their mode excitation mechanism only got explained 30 years ago by Moskalik & Dziembowski (1992), Gautschy & Saio (1993), and Dziembowski & Pamiatnykh (1993). The pulsations are driven by the heat mechanism (also known as $\kappa$ mechanism) active in the partial ionisation zone of iron-like elements in the outer envelope. However, while mode excitation in $\beta$ Cep stars is roughly understood, it remains incomplete. Indeed, more modes are detected in data than predicted by theory as revealed in some of the prototypical class members (Pamyatnykh et al., 2004; Daszyńska-Daszkiewicz & Walczak, 2010; Daszyńska-Daszkiewicz et al., 2013). Their observations show g-modes while they are predicted to be stable modes (i.e. not excited). Moravveji (2016) provided a summary of observed modes in $\beta$ Cep stars and illustrated the impact of higher opacities in the driving zone. Additional physical mechanisms may also be active in the interiors of such stars, while being absent in current excitation theory. Microscopic atomic diffusion, notably radiative levitation, was recently shown to be one such missing ingredient (Rehm et al., 2024).
@@
The interest in $\beta$ Cep stars came in waves of discoveries. The earliest work focused on the period determination from radial velocity time series of the very few known stars similar to $\beta$ Cep and the lack of understanding of the instability mechanism (e.g. review by Struve, 1952). In the following years, more $\beta$ Cep stars were discovered and the efforts to understand the pulsation mechanism were continued, although the physical origin was still elusive (Lesh & Aizenman, 1978). The discovery and explanation of the mode excitation mechanism (Pamyatnykh, 1999, for a summary) sparked new interest in $\beta$ Cep stars. Aerts et al. (1992, 1994) conducted successful mode identification from time series spectroscopy for a few bright class members, while Heynderickx et al. (1994) identified the degree of the dominant pulsation modes for a substantial number of class members from multi-colour photometry. While both methods worked, only a limited number of modes per star could be identified – often just one or two. Recently, Shitrit & Arcavi (2024) showed that ground-based multi-colour photometry is in principle still a viable way to achieve mode identification for asteroseismology of $\beta$ Cep stars, complementary to single-band modern space-based photometry.
@@
Inferences of a $\beta$ Cep star interior were achieved thanks to the 21 year-long photometric monitoring of HD 129929 (Aerts et al., 2003, 2004b; Dupret et al., 2004). These three dedicated studies kick-started $\beta$ Cep asteroseismology from the identification of six modes: a radial mode, a rotationally split p-mode triplet and two components of a g-mode quintuplet. It led to the first seismic probing of the internal rotation of a main-sequence star other than the Sun, with the bottom of the envelope rotating about 4 times faster than the outer envelope. Inspired by this result, in coordinated efforts months-long multi-site campaigns involving tens of astronomers worldwide were organised. As such, four of the brightest $\beta$ Cep stars got scrutinised seismically from combined multi-colour photometric and spectroscopic time-series data. These stars are still the best modelled $\beta$ Cep stars today, thanks to the unambiguous identification of their modes: $\nu\,$Eri (Aerts et al., 2004a; De Ridder et al., 2004; Pamyatnykh et al., 2004), $\theta\,$Oph (Handler et al., 2005; Briquet et al., 2005, 2007), 12 Lac (Handler et al., 2006; Dziembowski & Pamyatnykh, 2008; Desmet et al., 2009), and V2052 Oph (Handler et al., 2012; Briquet et al., 2012). The seismic estimates of the internal rotation, age (in terms of core hydrogen mass fraction), and core boundary mixing are summarised in Bowman (2020).
@@
With the new era of space photometry, asteroseismology has gained a treasure trove of high-quality, high-cadence data for many pulsating stars (Aerts, 2021; Kurtz, 2022, for recent reviews). For $\beta$ Cep stars in particular, the earliest space-based observations with the Microvariability and Oscillations of Stars (MOST) satellite revealed multi-periodic behaviour of $\delta$ Ceti, which was thought to be a mono-periodic radial pulsator until that time (Aerts et al., 2006). The CoRoT mission advanced our knowledge of only three more $\beta$ Cep stars (Belkacem et al., 2009; Degroote et al., 2009; Briquet et al., 2011; Aerts et al., 2019) and the Kepler space satellite did not reveal any additional $\beta$ Cep pulsators with identified modes (Balona et al., 2011; Lehmann et al., 2011). This led to a paucity of research on these stars with space-based time series photometry, although asteroseismology as a whole was revolutionised by the Kepler mission.
@@
Already with the first two sectors of the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2015) mission it became obvious that massive star asteroseismology entered a new era (Bowman et al., 2019; Pedersen et al., 2019; Burssens et al., 2020). Shi et al. (2024) provide a catalogue of all $\beta$ Cep stars observed by TESS with 2 min-cadence data. Southworth & Bowman (2022) identified eight $\beta$ Cep pulsators in eclipsing binaries, while Eze & Handler (2024) recently expanded this list to 78 systems. However, to best study these stars long-baseline time series observations are needed to resolve the anticipated close-by frequencies and identify the pulsation modes.
@@
Burssens et al. (2023) analysed the 352 d-long light curve of the bright $\beta$ Cep star HD 192575 located in the TESS Northern continuous viewing zone and were able to model its structure parameters with high precision as the fifth $\beta\,$Cep star with tight seismic constraints. They also estimated its internal rotation profile from forward modelling but encountered large uncertainty due to the use of only three rotationally split multiplets. Vanlaer et al. (in prep.) meanwhile analysed the 5-year TESS light curve (with twice a gap of about a year) and resolved some of the ambiguity in the mode identification, allowing the authors to perform a rotation inversion.
@@
Next to CoRoT, Kepler, and TESS, the Gaia mission has proven to be a highly successful gateway to discover pulsating stars. Despite not being designed for asteroseismology, the recurring visits to millions of stars in a quasi-random pattern allowed for the discovery of some 100,000 new main-sequence pulsators (Gaia Collaboration et al., 2023). Hey & Aerts (2024) revisited these stars and analysed their first two years of TESS photometry, confirming the main pulsation frequency and classification for more than 80% of these Gaia-identified pulsators. Given these findings, we are encouraged to focus on the 222 $\beta$ Cep stars in the sample of Gaia Collaboration et al. (2023) and to follow up on the work of Hey & Aerts (2024) by considering the available TESS data.
@@
Here, we use TESS light curves with a considerably longer time base than Hey & Aerts (2024), enabling as such maximal frequency resolution. Our aim is to hunt for resolved low-(radial) order pressure and gravity modes and rotational splitting in the 222 Gaia-discovered $\beta$ Cep stars, opening pathways to mode identification and seismic modelling from combined TESS photometry and Gaia data, in a homogeneous way for an as large as possible ensemble of $\beta$ Cep  stars with publicly available light curves assembled with both satellites.
@@ 2 Data reduction
We selected all 222 variable stars classified by Gaia Collaboration et al. (2023) as $\beta$ Cep stars as our initial sample. Since our work focuses on the synergy between the TESS and Gaia DR3 light curves, we only included these stars. Yet, we emphasise that the methods developed in this work could potentially be applied to TESS observations of known $\beta$ Cep stars with Gaia time series photometry. To prepare for the data reduction, we obtained for each star on our target list its position from the Gaia archive and cross-matched it with the TESS input catalogue (Stassun et al., 2019; Paegert et al., 2021) to obtain its TIC ID.
@@ 2.1 TESS data reduction and light curve detrending
We employed the TESS data reduction pipeline tessutils (Garcia et al., 2022) which enables the end-to-end creation of detrended light curves from TESS full-frame images (FFI) and is optimised for the asteroseismic analysis. We briefly recall its work-flow here. For every TIC number, tessutils downloads cut-outs from the FIIs using tesscut (Brasseur et al., 2019). In the present work we chose a cut-out size of 25 px by 25 px. Subsequently, tessutils searches for the optimal aperture for each sector while trying to keep the contamination from neighbouring stars below 1 %. Afterwards, it extracts the light curve with the chosen aperture.
Since many of our stars are located in crowded fields near the Galactic plane or in open clusters, we weakened the contamination criterion for stars with a large number of nearby sources. As the targeted $\beta$ Cep stars are typically among the brightest stars in the field they provide the main flux contribution. To reduce the influence of neighbouring stars, we set the aperture to a single pixel centred on the target’s position. With this approach, we were able to obtain light curves for 37 stars that would otherwise have (partially) been missed.
After aperture photometry, tessutils detrends the extracted light curves per sector with an automated principal component analysis with up to even components to remove long-term trends in the data. This is justified for the asteroseismic analysis of $\beta$ Cep stars as we are interested in variability with periodicities of typically several hours for the individual modes. The normalised light curves were binned to the longest cadence of the TESS photometry, that is 30 min. This guarantees a homogeneous treatment involving all TESS FFI data of the sample stars discovered from Gaia, while it leads to excellent agreement among the dominant frequencies detected in the TESS and Gaia DR3 light curves, as we will show below. We tested that relying on the higher-cadence TESS FFI data does not change the results for our $\beta$ Cep population study. Finally, the data of all sectors were stitched to a single long-term light curve per star.
Despite the initial detrending of the light curves, we still noticed significant instrumental trends in many light curves. Therefore, we manually selected and removed parts of the light curve where instrumental effects were dominant. An example of such a light curve is shown in Fig. 17. To remove the remaining non-astrophysical trends, we fitted and subtracted a second order polynomial to sections of the light curve separated by a gap in the data of at least three hours. In some cases, we manually selected the section to ensure a smooth detrending. Afterwards, we $\sigma$-clipped each section with a threshold of 5 $\sigma$. Finally, in a handful of exceptional cases, an entire TESS sector was removed from the light curve due to excessive background flux or noise.
Out of our initial sample of 222 stars, 216 have been observed by TESS and we were able to obtain time series photometry for all of them. The number of sectors in the light curves range from one to seven with a median of four visits by TESS over the current mission lifetime spanning about five years.
@@ 2.2 Frequency extraction from pre-whitening
The method of pre-whitening consists in identifying the dominant signal frequency in the Fourier transform of the light curve, finding the amplitude and phase of that peak frequency and subtracting the matching sinusoid from the light curve to construct a residual light curve. Subsequently, the next strongest frequency is identified in the residuals, a sinusoid is built, and subtracted, etc. This procedure is followed until a user-specified stop criterion is reached (see, e.g. Chapter 5 in Aerts et al., 2010, for details).
In the case of our sample of 216 $\beta$ Cep stars, each TESS light curve has its own complicated window function resulting from the month-long sectors of the TESS observations and the long gaps in between them. In order to extract all significant frequencies up to the Nyquist frequency of 24 d-1, we used a customised version of the pre-whitening module of the publicly available frequency analysis code STAR SHADOW (IJspeert et al., 2024). While the reader is referred to that paper for details on the methodology, we point out that STAR SHADOW’s pre-whitening procedure is computationally efficient because it does not rely on time-consuming non-linear regression for all extracted signals simultaneously. Rather, it only resorts to non-linear regression for those signals that strongly affect each other. On top of that, STAR SHADOW is well suited to analyse blended frequency peaks in Fourier spectra, which are quite common in our set of light curves due to the limited coverage of the baseline.
During the frequency extraction via pre-whitening, STAR SHADOW relies on the Bayesian Information criterion (BIC) as its stopping criterion (Schwarz, 1978). In order to avoid over-extraction of frequencies from the high-quality space photometry, we adopted $\Delta$BIC$>$10 as criterion in each pre-whitening step. This means that a new frequency is only accepted as present in the data at a particular stage of pre-whitening if the BIC of its residual curve is reduced by at least a value of 10 compared to the BIC of the light curve in the previous stage of the pre-whitening. This is more conservative than the standard cutoff value of 2 adopted in many statistical applications (Kass & Raftery, 1995).
We employed a hybrid method for the selection of the next signal to be pre-whitened. Herein, the pre-whitening initially uses the concept of amplitude hinting (e.g. Van Beeck et al., 2021) and swaps to signal-to-noise ratio (SNR) hinting if a frequency is initially rejected, where the noise level is computed over a range of 1 d-1 in the Fourier spectrum. If that or another potential signal is rejected again, the pre-whitening is terminated. For the final significance assessment of each signal, STAR SHADOW demands the signal’s SNR to be greater than the threshold defined in Eq. 6 of (Baran & Koen, 2021). We increased the SNR-threshold by 0.25 for light curves with long gaps as recommended by Baran & Koen (2021).

### v2

@@ Abstract
Context. The Gaia mission detected many new candidate $\beta$ Cephei ($\beta$ Cep) pulsators, whose variability classification has since been confirmed from Transiting Exoplanet Survey Satellite (TESS) space photometry of the nominal mission.
Aims. We aim to analyse all currently available TESS data for these $\beta$ Cep pulsators, of which 145 were new discoveries, in order to exploit their asteroseismic potential. Although they are of critical importance to improve evolution models of massive stars, $\beta$ Cep stars are under-represented in the current space photometry revolution.
Methods. We extracted light curves for 216 stars from the TESS full-frame images and performed frequency analysis by means of pre-whitening. Based on Gaia Data Release 3, we deduced stellar properties and compared them to those of known $\beta$ Cep stars from the literature. We developed a methodology to identify the dominant pulsation modes of the $\beta$ Cep stars from the detection of rotationally-split multiplets and Gaia and TESS amplitude ratios. We used grid modelling to gain insights into the population of $\beta$ Cep stars.
Results. Combining TESS and Gaia, we successfully identified the mode degrees for 148 stars in our sample. We find the majority to have a dominant dipole non-radial mode. Many non-radial modes show splittings in their TESS frequency spectra, which we use to calculate their envelope rotation, spin parameter, and the level of differential envelope-to-surface rotation. For the latter, we find an upper limit of about 3. We also provide relative frequency asymmetries within the multiplets, ranging from -0.15 to 0.15 with most being positive. Based on grid modelling, we provide mass, convective core mass, and age distributions for 119 stars.
Conclusions. Our sample enables asteroseismology of $\beta$ Cep pulsators as a population. Our study prepares for future detailed modelling based on individual frequencies of identified modes leading towards a better understanding of these massive pulsators.
@@ 1 Introduction
$\beta$ Cephei ($\beta$ Cep) stars are core-hydrogen burning massive stars ($8\lesssim M/M_{\sun}\lesssim 25$) pulsating in lower-order pressure (p-) or gravity (g-) modes. Despite being known for more than a century (Frost, 1902, 1906; Guthnick, 1913) their mode excitation mechanism was only explained 30 years ago by Moskalik & Dziembowski (1992), Gautschy & Saio (1993), and Dziembowski & Pamiatnykh (1993). The pulsations are driven by the heat mechanism (also known as $\kappa$ mechanism) active in the partial ionisation zone of iron-like elements in the outer envelope. However, while mode excitation in $\beta$ Cep stars is roughly understood, it remains incomplete. Indeed, more modes are detected in data than predicted by theory as revealed in some of the prototypical class members (Pamyatnykh et al., 2004; Daszyńska-Daszkiewicz & Walczak, 2010; Daszyńska-Daszkiewicz et al., 2013). Their observations show g-modes while they are predicted to be stable modes (i.e. not excited). Moravveji (2016) provided a summary of observed modes in $\beta$ Cep stars and illustrated the impact of higher opacities in the driving zone. Additional physical mechanisms may also be active in the interiors of such stars, while being absent in the current excitation theory. Microscopic atomic diffusion, notably radiative levitation, was recently shown to be one such missing ingredient (Rehm et al., 2024).
@@
The interest in $\beta$ Cep stars came in waves of discoveries. The earliest work focused on the period determination from radial velocity time series of the very few known stars similar to $\beta$ Cep [^1] Due to their differences in beating structure in the light and radial-velocity curves, some $\beta$ Cep stars were initially classified as $\beta$ Canis Majoris stars (van Hoof, 1957). Eventually, $\beta$ Cep stars became the accepted name of this class of variable stars. and the lack of understanding of the instability mechanism (e.g. review by Struve, 1952). In the following years, more $\beta$ Cep stars were discovered and the efforts to understand the pulsation mechanism were continued, although the physical origin was still elusive (Lesh & Aizenman, 1978). The discovery and explanation of the mode excitation mechanism (Pamyatnykh, 1999, for a summary) sparked new interest in $\beta$ Cep stars. Aerts et al. (1992, 1994) conducted successful mode identification from time series spectroscopy for a few bright class members, while Heynderickx et al. (1994) identified the degree of the dominant pulsation modes for a substantial number of class members from multi-colour photometry. While both methods worked, only a limited number of modes per star could be identified – often just one or two. Recently, Shitrit & Arcavi (2024) showed that ground-based multi-colour photometry is still a viable way to achieve mode identification for asteroseismology of $\beta$ Cep stars, complementary to single-band modern space-based photometry.
@@
Inferences of a $\beta$ Cep star interior were achieved thanks to the 21 year-long photometric monitoring of HD 129929 (Aerts et al., 2003, 2004b; Dupret et al., 2004). These three dedicated studies kick-started $\beta$ Cep asteroseismology from the identification of six modes: a radial mode, a rotationally split p-mode triplet and two components of a g-mode quintuplet. It led to the first seismic probing of the internal rotation of a main-sequence star other than the Sun, with the bottom of the envelope rotating about four-times faster than the outer envelope. Inspired by this result, in coordinated efforts months-long multi-site campaigns involving tens of astronomers worldwide were organised. As such, four of the brightest $\beta$ Cep stars were scrutinised seismically from combined multi-colour photometric and spectroscopic time-series data. These stars are still the best modelled $\beta$ Cep stars today, thanks to the unambiguous identification of their modes: $\nu\,$Eri (Aerts et al., 2004a; De Ridder et al., 2004; Pamyatnykh et al., 2004), $\theta\,$Oph (Handler et al., 2005; Briquet et al., 2005, 2007), 12 Lac (Handler et al., 2006; Dziembowski & Pamyatnykh, 2008; Desmet et al., 2009), and V2052 Oph (Handler et al., 2012; Briquet et al., 2012). The seismic estimates of the internal rotation, age (in terms of core hydrogen mass fraction), and core boundary mixing are summarised in Bowman (2020).
@@
With the new era of space photometry, asteroseismology has gained a treasure trove of high-quality, high-cadence data for many pulsating stars (see Aerts, 2021; Kurtz, 2022, for recent reviews). For $\beta$ Cep stars in particular, the earliest space-based observations with the Microvariability and Oscillations of Stars (MOST) satellite revealed multi-periodic behaviour of $\delta$ Ceti, which was thought to be a mono-periodic radial pulsator until then (Aerts et al., 2006). The CoRoT mission advanced our knowledge of only three more $\beta$ Cep stars (Belkacem et al., 2009; Degroote et al., 2009; Briquet et al., 2011; Aerts et al., 2019) and the Kepler space satellite did not reveal any additional $\beta$ Cep pulsators with identified modes (Balona et al., 2011; Lehmann et al., 2011). This led to a paucity of research on these stars with space-based time series photometry, although asteroseismology as a whole was revolutionised by the Kepler mission.
@@
Already with the first two sectors of the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2015) mission it became obvious that massive star asteroseismology entered a new era (Bowman et al., 2019; Pedersen et al., 2019; Burssens et al., 2020; Balona & Ozuyar, 2020; Sharma et al., 2022). Shi et al. (2024) provide a catalogue of all $\beta$ Cep stars observed by TESS with 2 min-cadence data. Southworth & Bowman (2022) identified eight $\beta$ Cep pulsators in eclipsing binaries, while Eze & Handler (2024) recently expanded this list to 78 systems. However, to best study these stars long-baseline time series observations are needed to resolve the anticipated close-by frequencies and identify the pulsation modes.
@@
Burssens et al. (2023) analysed the 352 d-long light curve of the bright $\beta$ Cep star HD 192575 located in the TESS Northern continuous viewing zone and were able to model its structure parameters with high precision as the fifth $\beta\,$Cep star with tight seismic constraints. They also estimated its internal rotation profile from forward modelling but encountered large uncertainty due to the use of only three rotationally split multiplets. Vanlaer et al. (in prep.) meanwhile analysed the 5-year TESS light curve (with two year-long gaps) and resolved some of the ambiguity in the mode identification, allowing the authors to perform a rotation inversion.
@@
Next to CoRoT, Kepler, and TESS, the Gaia mission has proven to be a highly successful gateway to discover pulsating stars. Despite not being designed for asteroseismology, the recurring visits to millions of stars in a quasi-random pattern allowed for the discovery of ¿100,000 new main-sequence pulsators (Gaia Collaboration et al., 2023). Hey & Aerts (2024) revisited these stars and analysed their first two years of TESS photometry, confirming the main pulsation frequency and classification for more than 80% of these Gaia-identified pulsators. Given these findings, we are encouraged to focus on the 222 $\beta$ Cep stars in the sample of Gaia Collaboration et al. (2023) and to follow up on the work of Hey & Aerts (2024) by considering all available TESS data.
@@
Here, we use TESS light curves with a considerably longer time base than Hey & Aerts (2024), thus enabling improved frequency resolution. Our aim is to search for resolved low-(radial) order pressure and gravity modes and rotational splittings in the 222 Gaia-detected $\beta$ Cep stars, opening pathways to mode identification and seismic modelling from combined TESS photometry and Gaia data, in a homogeneous way for an as large as possible ensemble of $\beta$ Cep  stars with publicly available light curves assembled with both satellites.
@@ 2 Data reduction
We selected all 222 variable stars classified by Gaia Collaboration et al. (2023) as $\beta$ Cep stars as our initial sample. Since our work focuses on the synergy between the TESS and Gaia DR3 light curves, we only included these stars. Yet, we emphasise that the methods developed in this work could potentially be applied to TESS observations of known $\beta$ Cep stars with Gaia time series photometry. To prepare for the data reduction, we obtained for each star on our target list its position from the Gaia archive and cross-matched it with the TESS input catalogue (Stassun et al., 2019; Paegert et al., 2021) to obtain its TIC ID. Out of our initial sample of 222 stars, 216 have been observed by TESS and we can obtain time series photometry for them.
@@ 2.1 TESS data reduction and light curve detrending
We employed the TESS data reduction pipeline tessutils (Garcia et al., 2022) which enables the end-to-end creation of detrended light curves from TESS full-frame images (FFI) and is optimised for asteroseismic analysis. We briefly recall its work-flow here. For every TIC number, tessutils downloads cut-outs from the FFIs using tesscut (Brasseur et al., 2019). In the present work we chose a cut-out size of 25 px by 25 px. Subsequently, tessutils searches for the optimal aperture for each sector while trying to keep the contamination from neighbouring stars below 1 %. Afterwards, it extracts the light curve of that TESS sector as the total flux within the chosen aperture. The number of sectors in the light curves range from one to seven with a median of four visits by TESS over the current mission lifetime spanning about five years.
Fifteen of our stars are located in very crowded fields near the Galactic plane or in open clusters. Since the targeted $\beta$ Cep stars are typically among the brightest stars in the field they provide the main flux contribution. To reduce the influence of neighbouring stars, we set the aperture for these stars to a single pixel centred on the target’s position (see Appendix A for an example). With this approach, we were able to obtain light curves for stars that would otherwise have (partially) been missed.
After aperture photometry, tessutils detrends the extracted light curves per sector with an automated principal component analysis with up to seven components to remove long-term trends in the data. This is justified for the asteroseismic analysis of $\beta$ Cep stars as we are interested in variability with periodicities of typically several hours for the individual modes. The normalised light curves were binned to the longest cadence of the TESS photometry, that is 30 min. This guarantees a homogeneous treatment of all TESS FFI data of the sample stars detected by Gaia, while it leads to excellent agreement among the dominant frequencies detected in the TESS and Gaia DR3 light curves, as we will show below. We tested that restricting to the higher-cadence TESS FFI data does not change the results for our $\beta$ Cep population study. Finally, the data of all sectors were stitched to a single long-term light curve per star. However, despite the initial PCA detrending, we still noticed significant instrumental trends in many light curves. Therefore, we removed those parts of the light curves where instrumental effects were dominant and detrended the remaining light curve. The full description of the process and an example are given in Appendix A.
@@ 2.2 Frequency extraction from pre-whitening
For the extraction of frequencies, we used a pre-whitening procedure. This method consists in identifying the dominant signal frequency in the Fourier transform of the light curve, finding the amplitude and phase of it and subtracting the matching sinusoid from the light curve to construct a residual light curve. Subsequently, the next strongest frequency is identified in the residuals, a sinusoid is built, and subtracted, etc. This procedure is followed until a user-specified stop criterion is reached (see, e.g. Chapter 5 in Aerts et al., 2010, for details).
In order to extract all significant frequencies up to the Nyquist frequency of 24 d-1, we used a customised version of the pre-whitening module of the publicly available frequency analysis code STAR SHADOW (IJspeert et al., 2024). While the reader is referred to that paper for details on the methodology, we point out that STAR SHADOW’s pre-whitening procedure is computationally efficient because it does not rely on time-consuming non-linear regression for all extracted signals simultaneously. Rather, it only resorts to non-linear regression for those signals that strongly affect each other. On top of that, STAR SHADOW is well suited to analyse blended frequency peaks in Fourier spectra, which are quite common in our set of light curves due to the limited coverage of the baseline.
In each pre-whitening step, STAR SHADOW demands the Bayesian Information criterion (BIC) (Schwarz, 1978) decreases by at least 2 to accept a signal as significant. In order to discourage over-extraction of frequencies from the high-quality space photometry, we increased STAR SHADOW’s BIC-threshold in each pre-whitening step to 10. Additionally, we employed a hybrid method for the selection of the next signal to be pre-whitened. Herein, the pre-whitening initially uses amplitude hinting (e.g. Van Beeck et al., 2021) and swaps to signal-to-noise ratio (SNR) hinting once a frequency is rejected, where the noise level is computed over a range of 1 d-1 in the Fourier spectrum. If another potential signal is rejected, the pre-whitening is terminated. For the final significance assessment of each signal, STAR SHADOW demands the signal’s SNR to be greater than the threshold defined in Eq. 6 of Baran & Koen (2021). We increased the SNR-threshold by 0.25 for light curves with long gaps as recommended by Baran & Koen (2021).